# DIRECT: Deep Active Learning under Imbalance and Label Noise

### v3

@@ Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods. In this work, we conduct the first study of active learning under both class imbalance and label noise. We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage classic active learning theory and methods to address issues such as batch labeling and tolerance towards label noise. We present extensive experiments on imbalanced datasets with and without label noise. Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
@@ 1 Introduction
  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32). DIRECT consistently outperforms existing baseline algorithms by saving more than $60\%$ annotation cost compared to the best existing algorithm, and more than $80\%$ annotation cost compared to random sampling.
@@ 2 Related Work
Class-Balanced Deep Active Learning Active earning strategies sequentially and adaptively choose examples for annotation. Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong and Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have been shown to perform among the top when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2024a). More sophisticated methods have been proposed to optimize chosen examples’ uncertainty (Gal et al., 2017; Ducoffe and Precioso, 2018; Beluch et al., 2018), diversity (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as label noises (Khosla et al., 2022) or class imbalance Kothawade et al. (2021); Zhang et al. (2022, 2024a).
@@
Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications. Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2024b). To this end, Kothawade et al. (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances. GALAXY(Zhang et al., 2022) constructs one-dimensional linear graphs and applies graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain. While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation. In addition, bisection procedures are generally not robust against label noises, a prevalent challenge in real world annotation tasks. Our algorithm DIRECT mitigates all of the above shortcomings of GALAXY while outperforming it even with synchronous labeling and no label noise, beating GALAXY in its own game. Lastly, we distinguish our work from Zhang et al. (2024b), where the paper studies the algorithm selection problem. Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
@@
Agnostic Active Learning for Label Noise Label noise for active learning has been primarily studied under the extensive literature on agnostic learning. We refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion. All of these works, beginning with the seminal works by Balcan et al. (2006); Dasgupta et al. (2007), follow a familiar paradigm of disagreement based learning. This involves maintaining a version space of promising hypotheses at each time and constructing a disagreement region of unlabeled examples. For any unlabeled example in the disagreement region, there exists two hypotheses in the version disagreeing on their predictions. An example then chosen for annotation by sampling from a informative sampling distribution computed over the disagreement region. Several approaches have been proposed for computing such sampling distributions, e.g. Jain and Jamieson (2019); Katz-Samuels et al. (2020, 2021); Huang et al. (2015). As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al. (2021). We remark that these algorithms tend to be overly pessimistic in training deep neural nets, and this paper hopes to close this gap.
@@
Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature. Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021). In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon. This is fundamentally different from our work, where our annotator may provide noisy labels. Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang and Chaudhuri, 2015; Chen et al., 2022). The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on such instances. In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today. Recently, Khosla et al. (2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise. Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
To our knowledge, no deep active learning literature has studied the scenario where both imbalance and label noise present. Yet, this setting is the most prevalent in real-world annotation applications.
@@ 4.1 An 1-D Reduction to Agnostic Active Learning
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}$ is defined as $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}$. Here, $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times. The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1). Namely, with ties broken similar to above, $j^{\star}=\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$.
@@
$\left(q_{(1)}^{k}\leq\mkern-2.0mu{\cdots}\leq q_{(N)}^{k}\mkern-1.0mu{:}\text{ sorted permutation of\,}\{\widehat{p}_{i}^{k}\}_{i=1}^{N}\right)\mkern-1.0mu{\land}\left(q_{(i)}^{k}\mkern-1.0mu{=}q_{(i+1)}^{k}\Rightarrow[\widetilde{p}(x_{i})]_{k}\geq[\widetilde{p}(x_{i+1})]_{k}\right).$ (2)
@@
        Sort margin scores $0=q_{(0)}^{k}\leq q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ based on equation (2).
@@
        Identify threshold for class $k$: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x^{k}_{(i)},y^{k}_{(i)}),q^{k}_{(i)}\}_{i=1}^{N})$.
@@
4.2 Algorithm
We are now ready to state our algorithm DIRECT as shown in Algorithm 1. Each round of DIRECT follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class. We spend half each round’s budget for both phases.
During the first phase, to identify the optimal separation threshold for all classes, we loop over each class $k$ and run a agnostic active learning procedure for the corresponding 1-D class-k v.s. rest reduction. Our agnostic active learning subroutine is formally outlined in Algorithm 2, VReduce. Our method corresponds to the fixed-budget ACED Katz-Samuels et al. (2021) algorithm, adopted to the threshold classifier scenario. To explain the intuition about ACED in this scenario, we first note that in the separable case with $\mathcal{L}(h_{j^{\star}})=0$, we could simply run a bisection procedure to learn $j_{k}^{\star}$, the optimal separation threshold for the $k$-th class. However, as $j^{\star}$ cannot perfectly separate the classes, ACED circumvents this by maintaining a version space of possible thresholds, namely the interval $[I,J]$. Statistically with high likelihood, the optimal separation threshold $j_{k}^{\star}$ lies in this version space of $[I,J]$. During each of the $m$ rounds of VReduce, $B_{\text{parallel}}$ samples are annotated and the version space’s length is shrank by a factor of $1/c$. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left. The second phase of DIRECT simply annotates examples closest to each optimal separation threshold, aiming to annotate a class-balanced and uncertain examples.
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels with sorted uncertainty $\{x_{(i)}^{k},y_{(i)}^{k},q_{(i)}^{k}\}_{i=1}^{N}$ (unlabeled $y_{(i)}^{k}$ hidden to learner).
  Initialize: Version space as the shortest segment of indices $[I,J]$ such that: for each labeled example $x_{(i)}$ with $i\leq I$, $y_{(i)}=k$, and for each labeled example $x_{(j)}$ with $j\geq J$, $y_{(j)}\neq k$.
  Initialize: Number of iterations $m\leftarrow\frac{b}{B_{\text{parallel}}}$. Shrinking factor $c\leftarrow\sqrt[m]{J-I}$
  for $t=1,...,m$ do
     Uniformly sample $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}^{k}$, …, $x_{(J)}^{k}$ for annotation and insert to $L$.
     For each $0\leq s\leq N$, compute ^$\widehat{\mathcal{L}}^{k}(s)=\sum_{r\leq s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{r>s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}=k\}$.
     Shrink version space by $\frac{1}{c}$: $I,J\leftarrow\operatorname*{arg\,min}_{i,j\in[I,J]:j-i=\frac{1}{c}(J-I)}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}$.
  end forReturn: Updated labeled set $L$.
Algorithm 2 VReduce: Version Space Reduction
@@ 5 Experiments
For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.1, showing its superior label-efficiency and ability to accommodate batch labeling. In Section 5.2, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels. Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises. While we highlight many results in this section, see Appendix E for complete results.
Experiment Setups. Our experiments utilize $10$ imbalanced datasets derived from popular computer vision datasets. For the ResNet experiments, we utilize imbalanced versions of CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) datasets. For the LabelBench experiments, we utilize the FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021) datasets. We refer the readers to Appendix D for more details on our experiment setups.
(a) Imbalanced CIFAR-10, three classes
(b) Imbalanced CIFAR-100, ten classes
(c) Imbalanced SVHN, two classes
(d) PathMNIST, two classes
(e) FMoW, Balanced Pool Accuracy
(f) Imbalanced SVHN, two classes, # of minority labeled examples
5.1 Experiments under Imbalance, without Label Noise
For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. We briefly distinguish the algorithms into two categories. In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples. The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty. For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results, including BADGE, Margin Sampling, CORESET and GALAXY.
As highlighted in Figures 1(a) and 4(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments. In Figures 1(c) and 4(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments. Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 4(c). In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 4(f)). While Zhang et al. (2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy. While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances. When too few examples are from minority class, the performance of the minority classes could be significantly hindered. When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class. Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.
(a) Imbalanced CIFAR-10, two classes, 10% label noise
@@ 5.1 Experiments under Imbalance, without Label Noise
Figure 5: Performance of DIRECT against baseline algorithms under $10\%$ label noise. (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework. Results are averaged over four trials and the shaded areas represent standard errors around the mean.

### v4

# Improved Algorithm for Deep Active Learning under Imbalance via Optimal Separation

@@ Abstract
Class imbalance severely impacts machine learning performance on minority classes in real-world applications. While various solutions exist, active learning offers a fundamental fix by strategically collecting balanced, informative labeled examples from abundant unlabeled data. We introduce DIRECT, an algorithm that identifies class separation boundaries and selects the most uncertain nearby examples for annotation. By reducing the problem to one-dimensional active learning, DIRECT leverages established theory to handle batch labeling and label noise – another common challenge in data annotation that particularly affects active learning methods. Our work presents the first comprehensive study of active learning under both class imbalance and label noise. Extensive experiments on imbalanced datasets show DIRECT reduces annotation costs by over 60% compared to state-of-the-art active learning methods and over 80% versus random sampling, while maintaining robustness to label noise.
Deep Learning, Active Labeling, Class Imbalance
@@ 1 Introduction
  • We conduct experiments across 12 dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32). DIRECT consistently outperforms existing baseline algorithms by saving more than $60\%$ annotation cost compared to the best existing algorithm, and more than $80\%$ annotation cost compared to random sampling.
@@ 2 Related Work
Class-Balanced Deep Active Learning Active learning strategies sequentially and adaptively choose examples for annotation. Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have been shown to perform among the top when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2024a). More sophisticated methods have been proposed to optimize chosen examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as label noises (Khosla et al., 2022) or class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2024a).
@@
Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications. Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2024b; Xie et al., 2024). To this end, Kothawade et al. (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances. GALAXY(Zhang et al., 2022) constructs one-dimensional linear graphs and applies graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain. While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation. In addition, bisection procedures are generally not robust against label noises, a prevalent challenge in real world annotation tasks. Our algorithm DIRECT mitigates all of the above shortcomings of GALAXY while outperforming it even with synchronous labeling and no label noise, beating GALAXY in its own game. Lastly, we distinguish our work from Zhang et al. (2024b), where the paper studies the algorithm selection problem. Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
@@
Agnostic Active Learning for Label Noise Label noise for active learning has been primarily studied under the extensive literature on agnostic learning. We refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion. All of these works, beginning with the seminal works by Balcan et al. (2006); Dasgupta et al. (2007), follow a familiar paradigm of disagreement based learning. This involves maintaining a version space of promising hypotheses at each time and constructing a disagreement region of unlabeled examples. For any unlabeled example in the disagreement region, there exists two hypotheses in the version disagreeing on their predictions. An example then chosen for annotation by sampling from a informative sampling distribution computed over the disagreement region. Several approaches have been proposed for computing such sampling distributions, e.g. Jain & Jamieson (2019); Katz-Samuels et al. (2020, 2021); Huang et al. (2015). As described in Section 4.3, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al. (2021). We remark that these algorithms tend to be overly pessimistic in training deep neural nets, and this paper hopes to close this gap.
@@
Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature. Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021). In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon. This is fundamentally different from our work, where our annotator may provide noisy labels. Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022). The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on such instances. In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today. Recently, Khosla et al. (2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise. Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels. To our knowledge, no deep active learning literature has studied the scenario where both imbalance and label noise present. Yet, this setting is the most prevalent in real-world annotation applications.
@@ 4.1 An 1-D Reduction to Agnostic Active Learning
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}$ is defined as $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}$. Here, $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times. The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to equation 4.1. Namely, with ties broken similar to above, $j^{\star}=\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$.
@@
$\left(q_{(1)}^{k}\leq\mkern-2.0mu{\cdots}\leq q_{(N)}^{k}\mkern-1.0mu{:}\text{ sorted permutation of\,}\{\widehat{p}_{i}^{k}\}_{i=1}^{N}\right)\mkern-1.0mu{\land}$
$\left(q_{(i)}^{k}\mkern-1.0mu{=}q_{(i+1)}^{k}\Rightarrow[\widetilde{p}(x_{i})]_{k}\geq[\widetilde{p}(x_{i+1})]_{k}\right).$ (2)
@@
4.2 One-Dimensional Agnostic Active Learning
The key insight of our approach is to leverage the well-established theory of agnostic active learning for threshold classifiers, which provides robust guarantees even under label noise. This allows us to robustly identify the optimal separation threshold $j^{\star}$ from the 1-D reduction above.
Problem Formulation and Noise Handling. In the agnostic setting, we have a sorted sequence $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ with corresponding (possibly noisy) binary labels $y_{(1)},y_{(2)},\ldots,y_{(N)}$. The fundamental challenge is that no hypothesis $h_{j}\in\mathcal{H}$ may achieve zero empirical loss $\mathcal{L}(h_{j})$ due to label noise or model misspecification.
To formalize how our algorithm handles label noise, we consider the underlying conditional probability function $P(y_{i}|x_{i})$ for each data example $x_{i}$. Through the dimensionality reduction $x_{i}\rightarrow q_{i}$ (where $q_{i}$ is the real-valued sigmoid/margin score), we obtain an ordered set of 1-dimensional features $\{q_{i}\}_{i=1}^{N}$. This mapping induces a distribution $P(y_{i}|q_{i})$ over the reduced space, which naturally encodes any label noise present in the annotations.
Our objective is to find the threshold classifier $h_{j^{\star}}$ from the set of 1-dimensional threshold classifiers $\{h_{j}\}$ that minimizes the probability of error with respect to $P(y_{i}|q_{i})$. Crucially, our agnostic approach makes no assumptions about the form of $P(y_{i}|x_{i})$ or the induced $P(y_{i}|q_{i})$, allowing the algorithm to handle arbitrary noise models without requiring prior knowledge of the noise distribution.
Version Space Reduction. The VReduce algorithm (Algorithm 1) implements a version space approach that maintains an interval $[I,J]$ representing plausible optimal thresholds. The key principle is that if we observe labeled examples $x_{(i)}$ with $i\leq I$ mostly belong to class $1$, and examples $x_{(j)}$ with $j\geq J$ mostly belong to class $2$, then the optimal threshold $j^{\star}$ with high likelihood will lie within $[I,J]$.
The algorithm proceeds iteratively by: (1) maintaining and updating the version space interval $[I,J]$ based on observed labels, (2) sampling unlabeled examples uniformly within this interval to maximize information gain, (3) shrinking the version space based on empirical loss estimates that account for the noisy observations, and (4) repeating until the labeling budget is exhausted.
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, sorted examples $\{x_{(i)}^{k},y_{(i)}^{k},q_{(i)}^{k}\}_{i=1}^{N}$ (note $y_{(i)}^{k}$ of unlabeled examples are hidden to the learner).
  Initialize: Version space $[I,J]$ as the shortest interval such that: $\forall i\leq I$ with $x_{(i)}\in L$: $y_{(i)}=k$, and $\forall j\geq J$ with $x_{(j)}\in L$: $y_{(j)}\neq k$.
  Number of iterations $m\leftarrow b/B_{\text{parallel}}$. Shrinking factor $c\leftarrow(J-I)^{1/m}$.
  for $t=1,...,m$ do
     Sample $B_{\text{parallel}}$ unlabeled examples uniformly from $\{x_{(I)}^{k},\ldots,x_{(J)}^{k}\}$ and add to $L$.
     Compute empirical loss: ^$\widehat{\mathcal{L}}^{k}(s)=\sum_{r\leq s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{r>s:x_{(r)}\in L}\mathbf{1}\{y_{(r)}=k\}$.
     Update version space:$[I,J]\leftarrow\operatorname*{arg\,min}_{[i,j]:j-i=(J-I)/c}\max\{\widehat{\mathcal{L}}^{k}(i),\widehat{\mathcal{L}}^{k}(j)\}$.
  end forReturn: Updated labeled set $L$.
Algorithm 1 VReduce: Version Space Reduction for Threshold Learning
Theoretical Guarantees. The VReduce algorithm inherits robust theoretical properties from the agnostic active learning literature, achieving near-minimax and instance-optimal sample complexity bounds (Dasgupta et al., 2007; Hanneke et al., 2014). Unlike bisection-based methods that fail under label corruption, the disagreement-based framework (Balcan et al., 2006) provides natural robustness with high probability guarantees regardless of the underlying data distribution or noise model. Building on the ACED framework (Katz-Samuels et al., 2021), our algorithm extends these classical guarantees to practical batch settings through the $B_{\text{parallel}}$ parameter while preserving all theoretical properties, making it suitable for real-world annotation scenarios with multiple parallel annotators.
4.3 Algorithm
We are now ready to state our algorithm DIRECT as shown in Algorithm 2. Each round of DIRECT follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class using the agnostic active learning approach described above. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class. We spend half each round’s budget for both phases.
@@ 4.3 Algorithm
        Sort margin scores $0=q_{(0)}^{k}\leq q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ based on equation 4.1.
@@
        Identify threshold for class $k$: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{x^{k}_{(i)},y^{k}_{(i)},q^{k}_{(i)}\}_{i=1}^{N})$.
@@
(a) Imbalanced CIFAR-10, three classes
(b) Imbalanced CIFAR-100, ten classes
(c) Imbalanced SVHN, two classes
(d) PathMNIST, two classes
(e) FMoW, Balanced Pool Accuracy
(f) Imbalanced SVHN, two classes, # of minority labeled examples
During the first phase, to identify the optimal separation threshold for all classes, we loop over each class $k$ and run the agnostic active learning procedure VReduce for the corresponding 1-D class-k v.s. rest reduction. The second phase of DIRECT simply annotates examples closest to each optimal separation threshold, aiming to annotate a class-balanced and uncertain examples.
@@ 5 Experiments
For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.1, showing its superior label-efficiency and ability to accommodate batch labeling. In Section 5.2, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels. Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises. While we highlight many results in this section, see Appendix E for complete results. Our implementation is publicly available at: https://github.com/EfficientTraining/LabelBench/blob/main/LabelBench/strategy/strategy_impl/direct.py
Experiment Setups. Our experiments utilize $14$ imbalanced datasets derived from popular computer vision datasets. For the ResNet experiments, we utilize imbalanced and/or long-tail versions of CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) datasets. For the LabelBench experiments, we utilize the FMoW (Christie et al., 2018), iWildcam (Beery et al., 2021), iNaturalist (Van Horn et al., 2018) and ImageNet-LT (Deng et al., 2009) datasets. We refer the readers to Appendix D for more details on our experiment setups.
(a) CIFAR-10LT, 10 classes, 15% label noise
@@
Figure 5: Performance of DIRECT against baseline algorithms under label noise. (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework. Results are averaged over four trials and the shaded areas represent standard errors around the mean.
@@
5.1 Experiments under Imbalance, without Label Noise
For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. We briefly distinguish the algorithms into two categories. In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples. The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty.
Figure 7: Performance of different active learning algorithm across different classes for the ImageNet-LT dataset.
For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results, including BADGE, Margin Sampling, CORESET and GALAXY. As highlighted in Figures 1(a) and 4(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments. In Figures 1(c) and 4(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments. Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 4(c). In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 4(f)). While Zhang et al. (2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy. While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances. When too few examples are from minority class, the performance of the minority classes could be significantly hindered. When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class. Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.